OpenStudy (lin.ivory):

A baseball team plays in he stadium that holds 54000 spectators. With the ticket price at 10 the average attendance has been 21000. When the price dropped to 7, the average attendance rose to 27000. Find the a) demand function b) ticket price with maximum revenue

OpenStudy (anonymous):

Is this finance/economics? You can find the demand function by taking the slope, assuming it is a linear function. 21000-27000/10-7 = 6000/3 = -2000 The slope is negative so for every dollar increase, your attendance decreases by 2000. Likewise, for every dollar decrease, your attendance increases by 2000. Since the demand curve must pass through both points, you can pick either one and write it in the y=mx+b format. (y-21000)=-2000(x-10) =y-21000=-2000x+20000 =y=41000-2000x or y-27000=-2000(x-7) =y-27000=-2000x+14000 =y=41000-2000x same result.

OpenStudy (anonymous):

y represents attendance and x represents ticket price, if that wasn't clear =)

OpenStudy (lin.ivory):

what if i want x to be the number of spectator? how do i plug that back into the equation?

OpenStudy (lin.ivory):

wait what am i saying!! :P

OpenStudy (lin.ivory):

but then how do i find maximum revenue of the ticket price?